OpenStudy (anonymous):

is there any way to find the value of simultaneous equations including Y AND X as variables only in 30 secinds?

OpenStudy (amistre64):

Can you give an example?

OpenStudy (anonymous):

are you saying that u wanna find out the solution (x &y) when 2 equations are given?

OpenStudy (anonymous):

35x+7y+9=0 22x+18y-5=0

OpenStudy (anonymous):

i want a method by which i can find x amd y with in no time ..

OpenStudy (anonymous):

30 seconds what?

OpenStudy (anonymous):

ok i got wait

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

are these separate questions or system of equations?

OpenStudy (anonymous):

its 1 question... its system of equations

OpenStudy (anonymous):

i want to find the value of x and y.. i knw a lot of methods.. but i wantthe most effective one .. which take less t ime

OpenStudy (anonymous):

well it depends on ur ability in solving problems........

OpenStudy (anonymous):

ok..how u will prefer to solve it

OpenStudy (anonymous):

i will compare equations thats too quick just concentrate......

OpenStudy (anonymous):

i have a technique.. with which u can solve them with in no time.. :)

OpenStudy (anonymous):

ohky tell;-)

OpenStudy (anonymous):

cheating?;-)?;-)

OpenStudy (anonymous):

if u solve them by generalizing them.. ex. ax+by+c=0 dx+ey+f=0 your final answer will be x = (bf-ey) / (ae-db) y = (af-dc) / (ae-db) if u will concentarte u wil knw how can we solve it just by watching the equations

OpenStudy (anonymous):

so in my case 35x+7y+9=0 22x+18y-5=0 x = (7*5 - 18*9) / (35*18 - 22*7) y = (35*-5 - 22*9) / (35*18 - 22*7)

OpenStudy (anonymous):

and i knw every 1 can do multipkication and division :) so u solved the the equations with out crushing ur mind :)

OpenStudy (anonymous):

its gud but i am not going 2 experience iy;-)

OpenStudy (anonymous):

its upon u :) r u a teacher? and from where?

OpenStudy (anonymous):

no i am student 4m uet pakistan

OpenStudy (anonymous):

i am a civil engineer.. studied from cyprus

OpenStudy (anonymous):

ok pakistani?

OpenStudy (anonymous):

yupp :) and proud to be :)

OpenStudy (anonymous):

proud 2 b pakistani;-)

OpenStudy (anonymous):

aap kon si engineering kar rhi hain?

OpenStudy (anonymous):

electronics